# The Bathtub of European AI Governance: Identifying Technical Sandboxes as the Micro-Foundation of Regulatory Learning

### v1

@@ 1. Introduction
The regulation of AI has become a defining policy challenge of the digital era. As AI systems permeate economic, social, and political domains, governments worldwide face the difficult task of balancing innovation with protection, flexibility with accountability, and global competitiveness with domestic values. Yet, despite a shared sense of urgency, national and regional approaches to AI governance diverge sharply, reflecting distinct legal traditions and philosophies of regulation.
The European Union (EU) stands out for its historically anticipatory and top-down regulatory tradition, epitomised by the AI Act (58). Unlike jurisdictions such as the United States, which rely primarily on sectoral rules, voluntary standards, and ex-post liability (Administration, 2025), or China, where governance is framed through industrial policy and targeted regulatory interventions in prominent domains (Chen and Xu, 2025), the EU has opted for a comprehensive, fundamental rights-based, and risk-oriented framework (Almada and Petit, 2025). The Artificial Intelligence Act (AI Act) imposes ex-ante obligations, especially for high-risk systems, spanning risk management, data governance, technical documentation, record-keeping, transparency, and human oversight.
@@
However, effective collaboration requires more than political will; it demands shared terminology and a technical infrastructure that renders the dialogue between regulators and practitioners smooth, traceable, and evidence-based. Yet, this technical infrastructure to support the collaborative regulatory mechanisms is often lacking (Genicot and Moraes, 2025). To bridge the current gap between legal expectations and technical operationalisation, this paper investigates the AI Technical Sandbox (AITS) as the necessary technical backbone capable of producing the data required for mutual regulatory learning.
To support this aim, this work offers three main contributions. First, we establish a theoretical model of the AI Act’s governance structure, drawing from social sciences and policy learning theory to map the complex network of actors into micro, meso, and macro levels of regulatory learning. Using this model, we identify the providers and deployers at the micro level as those actors responsible for evidence generation, which highlights their fundamental role and the need for their active involvement in regulatory learning processes. Second, drawing on this functional perspective, we identify three distinct scenarios (self-assessment, regulatory sandboxes, and conformity assessment by a notified body) where the AITS serves as the essential technical engine for generating the data driving regulatory learning. Third, we make suggestions for operationalising these environments by detailing the crucial infrastructural requirements needed to transform isolated testing data into regulatory signals at scale.
@@ 2.1. Regulatory Learning
The AI Act, taking a general definition of AI system (Article 3), casts a large net to ensure wide coverage of the regulations (European Commission, 2025b). Within this broad scope, modern generative AI systems pose a particular challenge due to their emergent — not explicitly programmed — behaviour arising from the complex interplay of their components, causing parallels to be drawn to established complex system sciences (Holtzman et al., 2025). Reflecting that reality, these considerations are extended to not only the development and analysis of AI systems, but also their governance (Kolt et al., 2025). To retain agility in its fundamental rights, risk-based approach in light of such unpredictable developments, the AI Act integrates into the New Legislative Framework (NLF) (Veale and Zuiderveen Borgesius, 2021), setting horizontal essential requirements and delegating technical specifications to harmonised standards, implementing and delegating acts for vertical governance (Leyden, 2025). However, a critical gap currently exists as the vertical, sector-specific standards required for practical implementation, to be developed by bodies such as European Committee for Standardisation (CEN) and European Committee for Electrotechnical Standardisation (CENELEC), have yet to be approved by the EU Commission (A. Leyden (2025); 63).
@@
At the same time, the individual micro-level learning process is unable to produce a stimulus large enough to incite learning at the macro level. Through micro-micro transitions, in which groups of individuals interact, learning becomes a collective process, providing the signal to prompt potential learning at the macro level. Overall, the model captures the continuous feedback loop between the macro and micro levels by way of the macro-micro, micro-micro, and micro-macro interactions; it provides a causal model of how both levels influence each other’s learning processes, with the micro level providing the foundations for evidence generation for learning at the macro level.
However, some models for policy and regulatory learning do not neatly fit Coleman’s bathtub, due to a missing element mediating between the macro and micro levels, micro-micro transitions not fully capturing the aggregation of learning evidence between levels (Dunlop and Radaelli, 2017), or the need for making the intermediate level in organisational policy learning explicit (Zaki, 2025). Publications aiming to supplement the analysis of learning dynamics often introduce or specify the meso level to capture these organisational processes that mediate interactions between micro-scale and macro-scale learning (Welch et al., 2025; Jepperson and Meyer, 2011; Moyson et al., 2017). The meso level consists precisely of this aggregation layer at which organisations contribute to and participate in collective learning, ultimately feeding the macro level.
@@ 2.2. Sandboxes
Although the AI Act does not explicitly refer to assessment sandboxes, Article 58(2)(a) mandates AI regulatory sandboxes to support tools and infrastructure for testing, benchmarking, assessment, and explainability. This places technical evaluation at the core of the AIRS mandate.
In practice, AIRS operate along a spectrum reflecting the maturity and needs of participants. Some deployments emphasize legal and procedural guidance, while others integrate technical testing capabilities, provided either in-house or through independent evaluators. This distinction is analytical rather than structural. Operationally, an AIRS functions as a single, flexible environment that can combine regulatory oversight and technical assessment as required.
Regulatory and technical sandboxes therefore form a continuum. Regulatory sandboxes deliver legal certainty, interpretative guidance, and supervised real-world testing; technical sandboxes supply the infrastructure for rigorous evaluation. Together, they enable compliance-by-design, reinforce regulatory learning, and support the deployment of trustworthy AI systems with clarity, evidence, and confidence.
3. AITS in Regulatory Learning Scenarios
The regulatory learning space defined by the AI Act has been investigated before in literature. Particular focus has been put on the role of the AIRS (Buscemi et al., 2025; Ahern, 2025a), or an analysis of the mechanisms for adaptation within the AI Act (Schrepel, 2025). In addition, Lewis et al. (Lewis et al., 2025) have mapped the learning space, providing an extensive inventory of the specific activities identified within the AI Act, including the bi-directional flow between these activities. This paper builds on and complements the existing work by providing an analytical framework for the learning space based on the extended bathtub model mentioned in Section 2.1, aiming at identifying socio-technical mechanisms in the regulatory network. As such, while overlap with previous works exists, we aim to help complete the picture of regulatory learning, particularly through the identification of the AITS as a consistent, reproducible assessment methodology capable of serving as the micro-level evidence generator.
In the following, the functional and institutional interpretations of the AI Act’s regulatory learning landscape and its involved actors are explored. These perspectives are then applied to conceptually anchor and situate the identified scenarios in which the AITS is the central technological tool enabling its role as evidence generator.
@@ 3.1. The AI Act’s Bathtub
The AI Act, through Articles 64 to 70, defines a multi-actor structure aiming at a distributed governance model through the decentralisation of tasks. It attributes a diverse set of roles in enforcement and reporting to the defined actors, giving rise to complex interactions between them, marking the AI Act as a distinctly collaborative governance framework (Cancela-Outeda, 2024). Novelli et al. (Novelli et al., 2025) have provided a hierarchical analysis of the governance structure, decomposing the actors and their interplay into corporate, national, and supranational levels. However, while comprehensive in an administrative sense, a purely structural analysis is insufficient for the purpose of analysing the regulatory learning space specifically as it does not properly model the causal flow from the macro-micro pressure to the eventual micro-macro learning. This insufficiency is especially revealed by a closer examination of the role of the AI Office (Article 64). As it resides within the Directorate-General for Communications Networks, Content and Technology (DG CONNECT), it is a part of the EU Commission and therefore lacks legal autonomy. Nevertheless, the AI Office is, among other responsibilities, tasked with enforcement of the regulation with respect to General Purpose AI (GPAI) models, supporting coordination of implementation of the regulation by member states, and being involved in standardisation processes, granting it a certain — albeit ambiguous — degree of operational autonomy (Novelli et al., 2025; Cancela-Outeda, 2024; Graux et al., 2025). This disconnect between the AI Office’s legal and operational autonomy marks an example of quasi-agencification, an ongoing trend within the EU governance landscape (Verhoest, 2018; European Parliament. Directorate General for Parliamentary Research Services., 2018; Novelli et al., 2025).
@@
Applying this functional lens to the actors in the AI Act’s learning landscape allows mapping them, based on their tasks in both enforcement and learning, to a three-level framework constituted of the micro, meso, and macro levels. This mapping, presenting the causal flow of the regulatory learning space defined by the EU AI Act between these three levels, is shown in Figure 1. At the source of the regulatory learning pipeline is the EU AI Act itself, providing the legislative framework exerting pressure on the lowest level by demanding technical compliance with the legislation. This responsibility of demonstrating technical compliance with legal requirements belongs to providers, deployers, and developers of AI systems, including large firms, Small-medium enterprises, start ups, but also public institutions or even members of civil society, should the system be openly available to third party scrutiny. Thus, they form the micro level of the framework; their activities, in the design and conformity assessment of their AI system, are what produce the micro-level evidence that can enable data-driven adaptation of the macro level, i.e.  through direct amendments of the AI Act, delegated acts, or implementing acts. As the European Commission is the sole actor with the legal power to exert these adaptations, the Commission represents the macro level.
@@
In contrast to the micro and macro levels, the meso level consists of those actors who are tasked with translation of the macro-micro transition on the one hand and evidence aggregation of the micro-macro transition on the other, effectively constituting the mediation filter between the macro and micro actors. Considering the meso actors in terms of translation, in the form of enforcement or guidance, and aggregation, creates four different functional categories that the actors can be grouped into. As the AI Act lays out multiple tasks for each actor, an exclusive categorisation of an actor into a single category is not possible; nevertheless, the categorisation provides clarity in discussing the role of each actor along the enforcement and the learning axes.
@@
The first of these categories is comprised of the Enforcers, actors delegated with enforcement of the legal requirements. This category includes, at a national level, the Competent Authorities, i.e., the Market Surveillance Authorities and Notifying Authorities, as well as Notified Bodies. Together, they are tasked with monitoring and assessing conformity of the micro level actors with the legal requirements laid out by the AI Act. It should be noted here that the NA are indirect enforcers as they only appoint and monitor the NB who perform the third party conformity assessment of AI systems. The group is completed by supranational bodies in the form of the AI Board and the AI Office. The former, composed of Member State representatives, assumes an enforcing role through the coordination of the national CA, whereas the latter is tasked with direct monitoring of GPAI systems and supporting the AI Board administratively as secretariat. Collectively, these actors exert the necessary top-down pressure on the micro level, ensuring that the theoretical obligations of the Act are met with actual technical compliance.
@@
The macro-micro transition does not solely consist of enforcement activities. Instead, some actors defined by the AI Act are tasked with supporting activities through guidance and help in translating the legislation into technical requirements. These actors, notably the AI Board, the AI Office, and MSA, constitute the Meso-micro advisor category. In addition, the Scientific Panel is mandated to provide external expertise to MSA at their request, thus playing a crucial role in e.g., AIRS, but also with supporting the AI Office in their tasks. As such, they primarily act in a meso-meso advisory role, but their meso-micro advisory function is to provide direct operational support to the micro level, specifically through the coordination of administrative practices and the facilitation of AIRS. By managing these sandboxes and ensuring consistent application of rules across Member States, they create a controlled environment where developers can experiment and receive immediate regulatory feedback.
@@
In the micro-macro transition, the majority of the meso-level actors serve as the crucial aggregator of the micro-level evidence, acting as a filter for the signals enabling policy adaptation. The first category consisting of these aggregators are the Meso-macro advisors. These actors receive the evidence produced by the micro level and the enforcers to then pass it on to the macro level, serving as the primary feedback loop for the EU Commission. This category includes the Advisory Forum, which aggregates stakeholder perspectives, as well as the AI Board and AI Office in their reporting capacity. By synthesizing raw data, such as incident reports, market trends, or AIRS outcomes, into structured opinions and recommendations, they provide the Commission with the necessary signal to trigger legislative learning and adaptation.
@@
Lastly, the Standardisers category groups those actors who translate the legislative provisions into concrete technical and regulatory specifications. The major actor in this category are the European Standardization Bodies, such as CEN, CENELEC, and European Telecommunications Standards Institute (ETSI). They develop standards, including those at the request by the EU Commission, that then become regulatory instruments in the form of harmonised standards and ultimately serve as the key instruments for sector-specific, vertical governance by presuming conformity with the AI Act. However, this functional category also extends to actors such as the AI Office when they develop codes of practice and implementation guidelines. Like technical standards, these instruments serve to bridge the gap between abstract legal commands and practical application, providing the detailed specifications necessary for the micro level to demonstrate conformity. While serving as intermediary between legislation and technical requirements, these instruments also become tools in the bottom-up feedback to assess the feasibility and potentially adapt the legislation based on the micro-level experience.
@@
Nevertheless, the bathtub model in Figure 1 does not address all interactions between the mapped actors for the sake of visual clarity. In fact, the categorisation of the meso actors according to the four outlined categories confirms the complexity of the AI Act’s collaborative governance approach. Taking into account the interplay of actors reinforces the complexity, as there is not only a multiplicity of roles per actor, but also a non-unique assignment of actors per role. Moreover, both Table 2 and Figure 1 reveal a certain conflict as to the hierarchy of the actors. In this sense, the explicit definition of tasks in the AI Act provides precedence of the AI Board and AI Office’s input to the macro level as opposed to, e.g., standardisation bodies. However, only the national NB are truly subordinate in hierarchy to the NA as they rely on their certification. This further highlights the need for a well-defined, coherent evidence generation at the micro level to enable interoperable analysis between the evidence aggregated by each actor at the meso level, ultimately constituting an essential prerequisite for regulatory learning.
@@
AI Office Meso Direct enforcement for GPAI. Supports AIRS; coordinates practices. Aggregates GPAI risk evidence. Develops Codes of Practice.
AI Board Meso Coordinates national CA. Issues guidance; advises on AIRS. Aggregates Member State data. /
@@
Standardization Bodies Meso / / Feeds feasibility insights. Develops harmonised standards.
CA: MSA Meso (Nat) Market surveillance; sanctions. Manages AIRS. Reports enforcement outcomes and AIRS data. /
@@
Providers / Deployers Micro Compliance by design. Participate in AIRS. Generate post-market evidence. Implement standards.

### v2

@@ 1. Introduction
The regulation of AI has become a defining policy challenge of the digital era. As AI systems permeate economic, social, and political domains, governments worldwide face the difficult task of balancing innovation with protection, flexibility with accountability, and global competitiveness with domestic values. Yet, despite a shared sense of urgency, national and regional approaches to AI governance diverge sharply, reflecting distinct legal traditions and philosophies of regulation. The European Union (EU) stands out for its historically anticipatory and top-down regulatory tradition, exemplified by the AI Act (62). Unlike jurisdictions such as the United States, which rely primarily on sectoral rules, voluntary standards, and ex-post liability (Administration, 2025), or China, where governance is framed through industrial policy and targeted regulatory interventions in prominent domains (Chen and Xu, 2025), the EU has opted for a comprehensive, fundamental rights-based, and risk-oriented framework (Almada and Petit, 2025). The Artificial Intelligence Act (AI Act) imposes ex-ante obligations, especially for high-risk systems, spanning risk management, data governance, technical documentation, record-keeping, transparency, and human oversight.
@@
However, effective collaboration requires more than political will; it demands shared terminology and a technical infrastructure that renders the dialogue between regulators and practitioners smooth, traceable, and evidence-based. Yet, this technical infrastructure to support the collaborative regulatory mechanisms is often lacking (Genicot and Moraes, 2025). To bridge the current gap between legal expectations and technical operationalisation, this paper investigates AITSes as the necessary technical backbone capable of producing the data required for mutual regulatory learning.
To support this aim, this work offers three main contributions. First, we establish a theoretical model of the AI Act’s governance structure, drawing from social sciences and policy learning theory to map the complex network of actors into micro, meso, and macro levels of regulatory learning. Using this model, we identify the providers and deployers at the micro level as those actors responsible for evidence generation, which highlights their fundamental role and the need for their active involvement in regulatory learning processes. Second, drawing on this functional perspective, we identify three distinct scenarios (self-assessment, regulatory sandboxes, and conformity assessment by a notified body) where AITSes serve as the essential technical engine for generating the data driving regulatory learning. Third, we make suggestions for operationalising these environments by detailing the crucial infrastructural requirements needed to transform isolated testing data into regulatory signals at scale.
@@ 2.1. Regulatory Learning
The AI Act, taking a general definition of AI system (Article 3), casts a large net to ensure wide coverage of the regulations (European Commission, 2025b). Within this broad scope, modern generative AI systems pose a particular challenge due to their emergent — not explicitly programmed — behaviour arising from the complex interplay of their components, causing parallels to be drawn to established complex system sciences (Holtzman et al., 2025). Reflecting that reality, these considerations are extended to not only the development and analysis of AI systems, but also their governance (Kolt et al., 2025). To retain agility in its fundamental rights, risk-based approach in light of such unpredictable developments, the AI Act integrates into the New Legislative Framework (NLF) (Veale and Zuiderveen Borgesius, 2021), setting horizontal essential requirements and delegating technical specifications to harmonised standards, implementing and delegating acts for vertical governance (Leyden, 2025). However, a critical gap currently exists as the vertical, sector-specific standards required for practical implementation, to be developed by bodies such as European Committee for Standardisation (CEN) and European Committee for Electrotechnical Standardisation (CENELEC), have yet to be approved by the EU Commission (A. Leyden (2025); 67).
@@
At the same time, the individual micro-level learning process is unable to produce a stimulus large enough to incite learning at the macro level. Through micro-micro transitions, in which groups of individuals interact, learning becomes a collective process, providing the signal to prompt potential learning at the macro level. Overall, the model captures the continuous feedback loop between the macro and micro levels by way of the macro-micro, micro-micro, and micro-macro interactions; it provides a causal model of how both levels influence each other’s learning processes, with the micro level providing the foundations for evidence generation for learning at the macro level. However, some models for policy and regulatory learning do not neatly fit Coleman’s bathtub, due to a missing element mediating between the macro and micro levels, micro-micro transitions not fully capturing the aggregation of learning evidence between levels (Dunlop and Radaelli, 2017), or the need for making the intermediate level in organisational policy learning explicit (Zaki, 2025). Publications aiming to supplement the analysis of learning dynamics often introduce or specify the meso level to capture these organisational processes that mediate interactions between micro-scale and macro-scale learning (Welch et al., 2025; Jepperson and Meyer, 2011; Moyson et al., 2017). The meso level consists precisely of this aggregation layer at which organisations contribute to and participate in collective learning, ultimately feeding the macro level.
@@ 2.2. Sandboxes
Although the AI Act does not explicitly refer to assessment sandboxes, Article 58(2)(a) mandates AIRSes to support tools and infrastructure for testing, benchmarking, assessment, and explainability. This places technical evaluation at the core of the AIRS mandate. In practice, AIRSes operate along a spectrum reflecting the maturity and needs of participants. Some deployments emphasise legal and procedural guidance, while others integrate technical testing capabilities, provided either in-house or through independent evaluators. This distinction is analytical rather than structural. Operationally, an AIRS functions as a single, flexible environment that can combine regulatory oversight and technical assessment as required. Regulatory and technical sandboxes therefore form a continuum. Regulatory sandboxes deliver legal certainty, interpretative guidance, and supervised real-world testing; technical sandboxes supply the infrastructure for rigorous evaluation. Together, they enable compliance-by-design, reinforce regulatory learning, and support the deployment of trustworthy AI systems with clarity, evidence, and confidence.
3. AITSes as Enablers of Regulatory Learning
The regulatory learning space defined by the AI Act has been investigated before in literature. Particular focus has been put on the role of AIRSes (Buscemi et al., 2025; Ahern, 2025a), or an analysis of the mechanisms for adaptation within the AI Act (Schrepel, 2025). In addition, Lewis et al. (Lewis et al., 2025) have mapped the learning space, providing an extensive inventory of the specific activities identified within the AI Act, including the bi-directional flow between these activities. This paper builds on and complements the existing work by providing an analytical framework for the learning space based on the extended bathtub model mentioned in Section 2.1, aiming at identifying socio-technical mechanisms in the regulatory network. As such, while overlap with previous works exists, we aim to help complete the picture of regulatory learning, particularly through the identification of AITSes as a consistent, reproducible assessment methodology capable of serving as the micro-level evidence generator. In the following, the functional and institutional interpretations of the AI Act’s regulatory learning landscape and its involved actors are explored. These perspectives are then applied to conceptually anchor and situate the identified scenarios in which AITSes are the central technological tool enabling their role as evidence generator.
@@ 3.1. The AI Act’s Bathtub
The AI Act, through Articles 64 to 70, defines a multi-actor structure aiming at a distributed governance model through the decentralisation of tasks. It attributes a diverse set of roles in enforcement and reporting to the defined actors, giving rise to complex interactions between them, marking the AI Act as a distinctly collaborative governance framework (Cancela-Outeda, 2024). Novelli et al. (Novelli et al., 2025) have provided a hierarchical analysis of the governance structure, decomposing the actors and their interplay into corporate, national, and supranational levels. While such structural analyses are administratively comprehensive, they are not, on their own, sufficient to characterise the regulatory learning space. In particular, they tend to remain descriptive of institutional arrangements and competences, without explicitly modelling the causal pathways through which macro-level regulatory signals shape micro-level practices, or how evidence produced at the micro level is subsequently consolidated and informs macro-level adaptation. As a result, the dynamics of bidirectional regulatory learning may remain implicit rather than analytically specified.
This observation can be illustrated by considering the role of the AI Office under Article 64. The AI Office is situated within the European Commission, under DG CONNECT, and does not constitute a legally independent body. This institutional placement primarily reflects considerations of coordination and policy coherence. At the same time, it highlights the analytical challenge of distinguishing between governance structures designed for oversight and coordination, and mechanisms explicitly oriented toward the systematic aggregation and circulation of implementation-level evidence within a regulatory learning framework. Nevertheless, the AI Office is entrusted with a range of responsibilities, including the enforcement of the regulation with respect to General Purpose AI (GPAI) models, support for coordination of implementation across Member States, and participation in standardisation processes. Taken together, these tasks suggest a degree of operational discretion in practice, even if this discretion does not amount to formal institutional independence. From an analytical perspective, this positioning can be understood as conferring a form of limited and context-dependent operational autonomy, whose scope and implications for regulatory learning remain to be further clarified in the literature (Novelli et al., 2025; Cancela-Outeda, 2024; Graux et al., 2025). This disconnect between the AI Office’s legal and operational autonomy marks an example of quasi-agencification, an ongoing trend within the EU governance landscape (Verhoest, 2018; European Parliament. Directorate General for Parliamentary Research Services., 2018; Novelli et al., 2025).
@@
Applying this functional lens to the actors in the AI Act’s learning landscape allows mapping them, based on their tasks in both enforcement and learning, to a three-level framework constituted of the micro, meso, and macro levels. This mapping, presenting the causal flow of the regulatory learning space defined by the EU AI Act between these three levels, is shown in Figure 1. At the source of the regulatory learning pipeline is the EU AI Act itself, providing the legislative framework that exerts pressure on the lowest level by demanding technical compliance with the legislation. This responsibility of demonstrating technical compliance with legal requirements belongs to providers, deployers, and developers of AI systems, including large firms, Small-medium enterprises, startups, but also public institutions or even members of civil society, should the system be openly available to third party scrutiny. Thus, they form the micro level of the framework; their activities, in the design and conformity assessment of their AI system, are what produce the micro-level evidence that can enable data-driven adaptation of the macro level, i.e.  through direct amendments of the AI Act, delegated acts, or implementing acts. As the European Commission is the sole actor with the legal power to exert these adaptations, the Commission represents the macro level.
@@
In contrast to the micro and macro levels, the meso level consists of those actors who are tasked with translation of the macro-micro transition on the one hand and evidence aggregation of the micro-macro transition on the other, effectively constituting the mediation filter between the macro and micro actors. Considering the meso actors in terms of translation, in the form of enforcement or guidance, and aggregation, creates four different functional categories that the actors can be grouped into. As the example of the AI Office shows, the AI Act lays out multiple tasks for each actor. Therefore, we consider an exclusive categorisation of an actor into a single category not possible; nevertheless, the categorisation aims at providing clarity in discussing the role of each actor along the enforcement and the learning axes.
@@
The first of these categories consists of the Enforcers, actors delegated with enforcement of the legal requirements. This category includes, at a national level, the Competent Authorities, i.e., the Market Surveillance Authorities and Notifying Authorities, as well as Notified Bodies. Together, they are tasked with monitoring and assessing conformity of the micro level actors with the legal requirements laid out by the AI Act. It should be noted that NA can be characterised as indirect enforcers, in the sense that their enforcement role is exercised primarily through the designation, appointment, and oversight of NB, which in turn carry out third-party conformity assessments of AI systems. The group is completed by supranational bodies in the form of the AI Board and the AI Office. The former, composed of Member State representatives, assumes an enforcing role through the coordination of the national CA, whereas the latter is tasked with direct monitoring of GPAI systems and supporting the AI Board administratively as secretariat. Collectively, these actors exert the necessary top-down pressure on the micro level, ensuring that the theoretical obligations of the Act are met with actual technical compliance.
@@
The macro-micro transition does not solely consist of enforcement activities. Instead, some actors defined by the AI Act are tasked with supporting activities through guidance and help in translating the legislation into technical requirements. These actors, notably the AI Board, the AI Office, and MSA, constitute the Meso-micro advisor category. In addition, the Scientific Panel is mandated to provide external expertise to MSA at their request, thus playing a crucial role in e.g., AIRSes, but also with supporting the AI Office in their tasks. As such, they primarily act in a meso-meso advisory role, but their meso-micro advisory function is to provide direct operational support to the micro level, specifically through the coordination of administrative practices and the facilitation of AIRSes. By managing these AIRSes and ensuring consistent application of rules across Member States, they create a controlled environment where developers can experiment and receive immediate regulatory feedback.
@@
In the micro-macro transition, the majority of the meso-level actors serve as the crucial aggregator of the micro-level evidence, acting as a filter for the signals enabling policy adaptation. The first category consisting of these aggregators are the Meso-macro advisors. These actors receive the evidence produced by the micro level and the enforcers to then pass it on to the macro level, serving as the primary feedback loop for the EU Commission. This category includes the Advisory Forum, which aggregates stakeholder perspectives, as well as the AI Board and AI Office in their reporting capacity. By synthesising raw data, such as incident reports, market trends, or AIRS outcomes, into structured opinions and recommendations, they provide the Commission with the necessary signal to trigger legislative learning and adaptation.
@@
Lastly, the Standardisers category groups those actors who translate the legislative provisions into concrete technical and regulatory specifications. The major actor in this category are the European Standardisation Bodies, such as CEN, CENELEC, and European Telecommunications Standards Institute (ETSI). They develop standards, including those at the request by the EU Commission, that then become regulatory instruments in the form of harmonised standards and ultimately serve as the key instruments for sector-specific, vertical governance by presuming conformity with the AI Act. However, this functional category also extends to actors such as the AI Office when they develop codes of practice and implementation guidelines. Like technical standards, these instruments serve to bridge the gap between abstract legal commands and practical application, providing the detailed specifications necessary for the micro level to demonstrate conformity. While serving as intermediary between legislation and technical requirements, these instruments also become tools in the bottom-up feedback to assess the feasibility and potentially adapt the legislation based on the micro-level experience.
@@
Nevertheless, the bathtub model in Figure 1 does not aim to represent all interactions between the mapped actors, in the interest of visual clarity. Even with this abstraction, the categorisation of meso-level actors into the four outlined categories illustrates the structural complexity of the AI Act’s collaborative governance approach. Considering the interplay of actors further reinforces this complexity, as individual actors may assume multiple roles, while specific roles may be shared across multiple actors. Both Table 2 and Figure 1 also suggest ambiguities regarding the relative positioning of actors within the governance structure.
In particular, while the AI Act explicitly assigns certain tasks that prioritise the input of the AI Board and the AI Office at the macro level, for example relative to standardisation bodies, these assignments do not amount to a uniformly hierarchical ordering across all actors. Notably, national NB constitute the clearest instance of formal subordination, as their operation is contingent on designation and oversight by the NA. Taken together, this observation underscores the importance of well-defined and coherent evidence generation at the micro level. Such coherence is a prerequisite for interoperable analysis across meso-level actors and, ultimately, for the effective aggregation of evidence necessary to support regulatory learning.
@@
AI Office Meso Direct enforcement for GPAI. Supports AIRSes; coordinates practices. Aggregates GPAI risk evidence. Develops Codes of Practice.
AI Board Meso Coordinates national CA. Issues guidance; advises on AIRSes. Aggregates Member State data. /
@@
Standardisation Bodies Meso / / Feeds feasibility insights. Develops harmonised standards.
CA: MSA Meso (Nat) Market surveillance; sanctions. Manages AIRSes. Reports enforcement outcomes and AIRS data. /
@@
Providers / Deployers Micro Compliance by design. Participate in AIRSes. Generate post-market evidence. Implement standards.